# Scalable Oversight via Partitioned Human Supervision

### v1

@@ Abstract
As artificial intelligence (AI) systems approach and surpass expert human performance across a broad range of tasks, obtaining high-quality human supervision for evaluation and training becomes increasingly challenging. Our focus is on tasks that require deep knowledge and skills of multiple domains. Unfortunately, even the best human experts are knowledgeable only in a single narrow area, and will not be able to evaluate the correctness of advanced AI systems on such superhuman tasks. However, based on their narrow expertise, humans may provide a weak signal, i.e., a complementary label indicating an option that is incorrect. For example, a cardiologist could state that “this is not related to cardiology,” even if they cannot identify the true disease. Based on this weak signal, we propose a scalable oversight framework that enables us to evaluate frontier AI systems without the need to prepare the ground truth. We derive an unbiased estimator of top-1 accuracy from complementary labels and quantify how many complementary labels are needed to match the variance of ordinary labels. We further introduce two estimators to combine scarce ordinary labels with abundant complementary labels. We provide finite-sample deviation guarantees for both complementary-only and the mixed estimators. Empirically, we show that we can evaluate the output of large language models without the ground truth, if we have complementary labels. We further show that we can train an AI system with such weak signals: we show how we can design an agentic AI system automatically that can perform better with this partitioned human supervision. Our code is available at https://github.com/R-Yin-217/Scalable-Oversight-via-Human-Partitioned-Supervision.
@@ 1 Introduction
As foundation models and artificial intelligence (AI) systems (OpenAI, 2025a; Anthropic, 2025; DeepSeek, 2025; DeepMind, 2025) approach and in some areas surpass expert human performance, supervision itself becomes a key bottleneck. Current alignment pipelines such as supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF; Ziegler et al. (2019); Stiennon et al. (2020); Ouyang et al. (2022)), or reinforcement learning from verifiable rewards (RLVR; Shao et al. (2024); Yu et al. (2025); Lambert et al. (2025)) presuppose that humans can reliably provide supervision or design verifiers for training and evaluation.
Yet for the superalignment regime (Bowman et al., 2022), future models will tackle problems whose solutions are too technical or too cross-disciplinary for any single human to verify comprehensively. When we cannot produce ground truth or prepare automated verifiers, how should we evaluate and train such systems?
@@
We begin from an observation about the expertise in high-skill domains in the human society (Wuchty et al., 2007). As tasks grow in difficulty, human experts specialize ever more narrowly: cardiology vs. oncology in medicine, sector-specific analysts in finance, or sub-subfields in science. Specialists can often make local determinations with high precision, e.g., a cardiologist might say, “this is not a cardiology case,” while an oncologist might point out, “this pattern is never seen in the field of oncology.” While such judgments may fail to positively certify correctness, they readily certify that certain options are wrong. In other words, even when positive ground truth is out of reach, human experts can often provide reliable complementary labels (Ishida et al., 2017; Yu et al., 2018; Wang et al., 2024a; 2025): labels indicating one or more options that are definitely the incorrect.
@@
This paper develops a scalable oversight (Amodei et al., 2016) framework that treats such partitioned human supervision as a first-class supervision primitive. Our starting point is a multi-choice evaluation setting where ordinary (true) labels may be scarce or unavailable, but complementary labels arise naturally from partitioned expertise and are abundant. We show that one can utilize these complementary labels to propose an unbiased estimator of top-1 accuracy for any system under test, and further combine them with scarce ordinary labels for further refinement of the estimator. The resulting estimators allow us to (i) evaluate frontier models without preparing any ground truth, and (ii) use these evaluations as learning signals to design agentic systems on tasks beyond expert human abilities.
Figure 1: Proposed framework. Tasks beyond human capabilities are routed to a randomly selected domain expert indicates whether the query belongs to their field. A “no” response yields a complementary label, which is then used to evaluate or train an AI system.
@@
Our work complements many other active lines of research on scalable oversight. For example, weak-to-strong generalization (Burns et al., 2024) studies whether strong models trained on weak (noisy) supervisors can exceed those supervisors, possibly by eliciting knowledge that the strong model already possesses. Easy-to-hard generalization (Sun et al., 2024) trains evaluators on easier problems and applies them to supervise harder ones, by leveraging the assumption that evaluation is easier than generation. Debate (Khan et al., 2024) elicits truth via adversarial arguments that a judge selects among. Reinforcement learning from AI feedback or Constitutional AI (Bai et al., 2022) replaces human preferences with AI‑generated feedback.
Recursive task decomposition (Wu et al., 2021) breaks the problem into leaf subtasks that humans can directly label or compare, for tasks that are easily decomposable. We address a new, orthogonal axis: when human experts cannot supply correct answers for hard, cross-domain instances, can we use their partitioned “this option is wrong” judgments to obtain performance estimates and useful training signals?
@@ 2 Methodology
That is, for each item, the revealed wrong index is sampled uniformly from the K1K-1 non-$y$ classes. For an item with a complementary label $\bar{y}\neq y$, define $W=\mathbb{I}\{\hat{y}\neq\bar{y}\}\in\{0,1\}$. The assumption in Eq. 1 is enforced by the following data collection protocol. We assume that we have a set of questions $\{x_{i}\}^{N}_{i=1}$ and $K$ annotators, where each annotator is a human expert in the field of one of the classes. For each question $x_{i}$, choose an annotator uniformly at random (let’s say an annotator in charge of the $k$-th class is chosen), and ask “Is the answer of question $x_{i}$ class kk?”. If the answer is yes, we save this in the ordinary-label set as an item with (ordinary) class $k$ label. If the answer is no, we save this in the complementary-label set as an item with (complementary) class $k$ label.
Finally, we will have two disjoint evaluation sets, drawn from the same distribution and evaluated by the same system: 1) An ordinary-label set of size $n_{\mathrm{o}}$, yielding indicators $Z_{1},\ldots,Z_{n_{\mathrm{o}}}$ and total correct count $S_{\mathrm{o}}=\sum_{j=1}^{n_{\mathrm{o}}}Z_{j}$. 2) complementary-label set of size $n_{\mathrm{c}}$, yielding indicators $W_{1},\ldots,W_{n_{\mathrm{c}}}$ and total “weakly-correct” [^1] Here “weakly-correct” means that the model prediction is not equal to the complementary label. For example, when $K=4$, avoiding the complementary label still leaves a $\tfrac{2}{3}$ chance of being wrong. count $S_{\mathrm{c}}=\sum_{i=1}^{n_{\mathrm{c}}}W_{i}$. The total number of samples is $N=n_{\mathrm{o}}+n_{\mathrm{c}}$.
@@ 2.1 Estimator from Complementary Labels
Proof.
For a single item, under Eq. (1),
$=\Pr(\hat{y}=y)\cdot 1+\Pr(\hat{y}\neq y)\cdot\Big(1-\tfrac{1}{K-1}\Big)=\frac{A+K-2}{K-1},$
where $\mathbb{E}[\cdot]$ denote expectation. If we consider $\hat{q}=\tfrac{1}{n_{\mathrm{c}}}\sum_{i=1}^{n_{\mathrm{c}}}W_{i}$, an empirical version of $\mathbb{E}[W]$, this $\hat{q}$ is an unbiased estimator of $\mathbb{E}[W]$. Hence, $\widehat{A}_{\mathrm{comp}}=(K-1)\hat{q}-(K-2)$ is an unbiased estimator of $A$. ∎
This linear correction is precisely the 011 loss specialization of the risk-rewrite identity of Eqs. (8) and (10) in Ishida et al. (2019); replacing expectations by sample means yields our unbiased estimator. A full derivation is provided in Appendix A.
@@
$\mathrm{Var}(\widehat{A}_{\mathrm{comp}})=(K-1)^{2}\,\mathrm{Var}\!\left(\tfrac{1}{n_{\mathrm{c}}}\sum_{i=1}^{n_{\mathrm{c}}}W_{i}\right)=\frac{(K-1)^{2}}{n_{\mathrm{c}}}\,q(1-q)=\frac{(A+K-2)(1-A)}{n_{\mathrm{c}}},$ (3)
where $\mathrm{Var}(\cdot)$ denote variance. This expression depends on the unknown $A$. In practice, a plug-in estimator is used.
For the ordinary-label dataset, the number of correct predictions satisfies $S_{\mathrm{o}}\sim\mathrm{Bin}(n_{\mathrm{o}},A)$, where $\mathrm{Bin}(n,p)$ denotes the binomial distribution with parameters $n$ (number of independent trials) and $p$ (success probability in each trial). Thus, $\mathbb{E}\!\left[{S_{\mathrm{o}}}/{n_{\mathrm{o}}}\right]=A$, and $\mathrm{Var}\!\left({S_{\mathrm{o}}}/{n_{\mathrm{o}}}\right)={A(1-A)}/{n_{\mathrm{o}}}$, taken with respect to the underlying population distribution $\mathcal{D}$ over $(x,y)$. Comparing this with Eq. (3), the variance ratio is
@@ 2.2 Combining Ordinary and Complementary Labels
The corresponding minimum variance equals $\mathrm{Var}_{\min}(\widehat{A}_{\mathrm{mix}})=\frac{\tfrac{A(1-A)}{n_{\mathrm{o}}}\cdot\tfrac{(K-1)^{2}}{n_{\mathrm{c}}}q(1-q)}{\tfrac{A(1-A)}{n_{\mathrm{o}}}+\tfrac{(K-1)^{2}}{n_{\mathrm{c}}}q(1-q)}.$
Since $A$ is unknown, we use plug-in estimators: $\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{ord}})=\tfrac{\widehat{A}_{\mathrm{ord}}(1-\widehat{A}_{\mathrm{ord}})}{n_{\mathrm{o}}}$, and $\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{comp}})=\tfrac{(K-1)^{2}}{n_{\mathrm{c}}}\,\hat{q}(1-\hat{q}),\ \hat{q}=\tfrac{S_{\mathrm{c}}}{n_{\mathrm{c}}}.$ Then
@@
$\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{IVW}})=\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{ord}})\,\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{comp}})/(\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{ord}})+\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{comp}}))$.
Substituting $q=\tfrac{A+K-2}{K-1}$ into Eq. (6) gives
@@
If $n_{\mathrm{c}}=0$, Eq. (9) reduces to $\widehat{A}_{\mathrm{ord}}$. If $n_{\mathrm{o}}=0$, it reduces to $\widehat{A}_{\mathrm{comp}}$. For $K=4$, one always has $\hat{q}\in[\tfrac{2}{3},1]$ when $A\in[0,1]$, which bounds the complementary-component variance.
Moreover, a one-step Newton update around the truth shows that the IVW estimator with optimal fixed weights coincides with the ML estimator up to the first order [^2] A detailed proof is provided in Appendix C., and both are statistically efficient.
@@ 2.3 Theoretical analysis
We analyze the deviation of the complementary-label estimator $\widehat{A}_{\mathrm{comp}}$ under the uniform wrong-index assumption Eq. (1). Although $\widehat{A}_{\mathrm{comp}}$ is unbiased, its variance can be non-negligible at small sample sizes. We present two finite-sample deviation bounds: a distribution-free Hoeffding bound (Boucheron et al., 2003) and a data-dependent Bernstein bound (Maurer & Pontil, 2009) leveraging empirical variance. We then extend these to a fixed-weight mixture estimator.
2.3.1 Hoeffding and empirical Bernstein bounds for $\widehat{A}_{\mathrm{comp}}$
We derive a unified bound for $\widehat{A}_{\mathrm{comp}}$ by taking the minimum of the Hoeffding and empirical Bernstein inequalities. This formulation is both variance-free (via Hoeffding) and variance-adaptive (via Bernstein).
@@ 2.3.1 Hoeffding and empirical Bernstein bounds for $\widehat{A}_{\mathrm{comp}}$
Empirical Bernstein part. Apply the two-sided empirical Bernstein inequality (Maurer & Pontil, 2009) to $\hat{q}=\tfrac{1}{n_{\mathrm{c}}}\sum_{i=1}^{n_{\mathrm{c}}}W_{i}$ with $W_{i}\in[0,1]$:
[…]
2.3.2 Mixture estimator bounds via union bound
@@
We now state a Bernstein-type bound for the mixed estimator that complements the union-bound inequalities established in Proposition 3; its proof is deferred to Appendix D.
@@
This bound tightens Hoeffding’s inequality by incorporating both the variance term $v$ and the (range-driven) linear term $c$. For practical plug-in forms and further discussion, see Appendix E.
Formally, Eq. (10) is valid for any fixed choice of the weight $w\in[0,1]$ (e.g., a constant choice or an estimate obtained from an independent pilot split). If instead $w$ is chosen adaptively from the same evaluation data (e.g., the plug-in $\widehat{w}_{\mathrm{IVW}}$), the fixed-weight assumption is violated and the bound is no longer guaranteed. In such cases one can (i) revert to the union-bound inequalities (Proposition 3), which hold for arbitrary (possibly data-dependent) $w$ via a triangle inequality; (ii) restore the fixed-$w$ condition through sample splitting or cross-fitting (Chernozhukov et al., 2018; Wager & Athey, 2018); or (iii) restrict $w$ to a finite grid $G\subset[0,1]$ and apply a union bound with $\delta$ adjusted by $|G|$ (Boucheron et al., 2013; Shalev-Shwartz & Ben-David, 2014).

### v2

# Towards Scalable Oversight via Partitioned Human Supervision

@@ Abstract
As artificial intelligence (AI) systems approach and surpass expert human performance across a broad range of tasks, obtaining high-quality human supervision for evaluation and training becomes increasingly challenging. Our focus is on tasks that require deep knowledge and skills of multiple domains, where this bottleneck is severe. Unfortunately, even the best human experts are knowledgeable only in a single narrow area, and will not be able to evaluate the correctness of advanced AI systems on such superhuman tasks. However, based on their narrow expertise, humans may provide a weak signal, i.e., a complementary label indicating an option that is incorrect. For example, a cardiologist could state that “this is not related to any cardiovascular disease,” even if they cannot identify the true disease. Based on this weak signal, we propose a scalable oversight framework that enables us to evaluate frontier AI systems without the need to prepare the ground truth. We derive an unbiased estimator of top-1 accuracy from complementary labels and quantify how many complementary labels are needed to match the variance of ordinary labels. We further introduce two estimators to combine scarce ordinary labels with abundant complementary labels. We provide finite-sample deviation guarantees for both complementary-only and the mixed estimators. Empirically, we show that we can evaluate the output of large language models without the ground truth, if we have complementary labels. We further show that we can train an AI system with such weak signals: we show how we can design an agentic AI system automatically that can improve itself with this partitioned human supervision. Our code is available at Towards Scalable Oversight via Partitioned Human Supervision.
@@ 1 Introduction
As foundation models and artificial intelligence (AI) systems (OpenAI, 2025b; Anthropic, 2025; DeepSeek, 2025; DeepMind, 2025) approach and in some areas surpass expert human performance, supervision itself becomes a key bottleneck. Current alignment pipelines such as supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF; Ziegler et al. (2019); Stiennon et al. (2020); Ouyang et al. (2022)), or reinforcement learning from verifiable rewards (RLVR; Shao et al. (2024); Yu et al. (2025); Lambert et al. (2025)) presuppose that humans can reliably provide supervision or design verifiers for training and evaluation. Yet for the superalignment regime (Bowman et al., 2022), future models will tackle problems whose solutions are too technical or too cross-disciplinary for any single human to verify comprehensively. When we cannot produce ground truth or prepare automated verifiers, how should we evaluate and train such systems?
@@
We begin from an observation about the expertise in high-skill domains in the human society (Wuchty et al., 2007). As tasks grow in difficulty, human experts specialize ever more narrowly: cardiology vs. oncology in medicine, sector-specific analysts in finance, or sub-subfields in science. Specialists can often make local determinations with high precision, e.g., a cardiologist might say, “this is not a cardiovascular disease,” while an oncologist might point out, “this pattern is never seen in the field of oncology.” While such judgments may fail to positively certify correctness, they readily certify that certain options are wrong. In other words, even when positive ground truth is out of reach, human experts can often provide reliable complementary labels (Ishida et al., 2017; Yu et al., 2018; Wang et al., 2024a; 2025): labels indicating one or more options that are definitely the incorrect.
@@
This paper develops a scalable oversight (Amodei et al., 2016) framework that treats such partitioned human supervision as a first-class supervision primitive. Our starting point is a multi-choice evaluation setting where ordinary (true) labels may be scarce or unavailable, but complementary labels arise naturally from partitioned expertise and are abundant. We show that one can utilize these complementary labels to propose an unbiased estimator of top-1 accuracy for an AI system under test, and combine them with scarce ordinary labels for refinement of the estimator. The estimators allow us to (i) evaluate frontier models without preparing any ground truth, and (ii) use these evaluations as learning signals to design agentic systems on tasks beyond expert human abilities.
Figure 1: Superhuman tasks are routed to a human expert who can only give a weak signal that they are not capable of solving the task. We show that this weak signal, i.e., complementary label, can be used to evaluate/train an AI system.
@@
Our work complements many other active lines of research on scalable oversight. For example, weak-to-strong generalization (Burns et al., 2024) studies whether strong models trained on weak (noisy) supervisors can exceed those supervisors, possibly by eliciting knowledge that the strong model already possesses. Easy-to-hard generalization (Sun et al., 2024) trains evaluators on easier problems and applies them to supervise harder ones, by leveraging the assumption that evaluation is easier than generation. Debate (Khan et al., 2024) elicits truth via adversarial arguments that a judge selects among. Reinforcement learning from AI feedback or Constitutional AI (Bai et al., 2022) replaces human preferences with AI‑generated feedback. Recursive task decomposition (Wu et al., 2021) breaks the problem into leaf subtasks that humans can directly label or compare, for tasks that are easily decomposable. We address a new, orthogonal axis: when human experts cannot supply correct answers for hard, cross-domain instances, can we use their partitioned “this option is wrong” judgments to obtain performance estimates and useful training signals?
@@ 2 Methodology
That is, for each item, the revealed wrong index is sampled uniformly from the K1K-1 non-$Y$ classes. For an item with a complementary label $\bar{Y}\neq Y$, define $W=\mathbb{I}\{\hat{Y}\neq\bar{Y}\}\in\{0,1\}$. The assumption in Eq. 1 is enforced by the following data collection protocol. We assume that we have a set of questions $\{x_{i}\}^{N}_{i=1}$ and $K$ annotators, where each annotator is a human expert in the field of one of the classes. For each question $x_{i}$, choose an annotator randomly (let’s say an annotator in charge of the $k$-th class is chosen), and ask “Is the answer of question $x_{i}$ class kk?”. If the answer is yes, we save this in the ordinary-label set as an item with (ordinary) class $k$ label. If the answer is no, we save this in the complementary-label set as an item with (complementary) class $k$ label. Under this protocol, the uniform sampling assumption in Eq. 1 holds by construction (Appendix 1). Since each expert is queried with equal probability and only performs a binary compatibility check for their designated class, the potential for systematic annotation bias is substantially reduced.
Finally, we will have two disjoint evaluation sets, drawn from the same distribution and evaluated by the same system: 1) An ordinary-label set of size $n_{\mathrm{o}}$, yielding realizations $z_{1},\ldots,z_{n_{\mathrm{o}}}$ and total correct count $S_{\mathrm{o}}=\sum_{j=1}^{n_{\mathrm{o}}}z_{j}$. 2) complementary-label set of size $n_{\mathrm{c}}$, yielding realizations $w_{1},\ldots,w_{n_{\mathrm{c}}}$ and total “weakly-correct” [^1] Here “weakly-correct” means that the model prediction is not equal to the complementary label. For example, when $K=4$, avoiding the complementary label still leaves a $\tfrac{2}{3}$ chance of being wrong. count $S_{\mathrm{c}}=\sum_{i=1}^{n_{\mathrm{c}}}w_{i}$. The total number of samples is $N=n_{\mathrm{o}}+n_{\mathrm{c}}$.
@@ 2.1 Estimator from Complementary Labels
We show the proof in Appendix A. This linear correction is precisely the 011 loss specialization of the risk-rewrite identity of Eqs. (8) and (10) in Ishida et al. (2019); replacing expectations by sample means yields our unbiased estimator. A full derivation is provided in Appendix B.
@@
$\mathrm{Var}(\widehat{A}_{\mathrm{comp}})=(K-1)^{2}\,\mathrm{Var}\!\left(\tfrac{1}{n_{\mathrm{c}}}\sum_{i=1}^{n_{\mathrm{c}}}W_{i}\right)=\frac{(K-1)^{2}}{n_{\mathrm{c}}}\,q(1-q)=\frac{(A+K-2)(1-A)}{n_{\mathrm{c}}}.$ (3)
This expression depends on the unknown $A$. In practice, a plug-in estimator is used. For the ordinary-label dataset, the number of correct predictions satisfies $S_{\mathrm{o}}\sim\mathrm{Bin}(n_{\mathrm{o}},A)$, where $\mathrm{Bin}(n,p)$ denotes the binomial distribution with parameters $n$ (number of independent trials) and $p$ (success probability in each trial). Thus, $\mathbb{E}\!\left[{S_{\mathrm{o}}}/{n_{\mathrm{o}}}\right]=A$, and $\mathrm{Var}\!\left({S_{\mathrm{o}}}/{n_{\mathrm{o}}}\right)={A(1-A)}/{n_{\mathrm{o}}}$, where $\mathbb{E}[\cdot]$ and $\mathrm{Var}(\cdot)$ denote expectation and variance, respectively, taken with respect to the underlying population distribution $\mathcal{D}$ over $(x,y)$. Comparing this with Eq. (3), the variance ratio is
@@ 2.2 Combining Ordinary and Complementary Labels
The corresponding minimum variance equals $\mathrm{Var}_{\min}(\widehat{A}_{\mathrm{mix}})=\frac{\tfrac{A(1-A)}{n_{\mathrm{o}}}\cdot\tfrac{(K-1)^{2}}{n_{\mathrm{c}}}q(1-q)}{\tfrac{A(1-A)}{n_{\mathrm{o}}}+\tfrac{(K-1)^{2}}{n_{\mathrm{c}}}q(1-q)}.$ Since $A$ is unknown, we use plug-in estimators: $\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{ord}})=\tfrac{\widehat{A}_{\mathrm{ord}}(1-\widehat{A}_{\mathrm{ord}})}{n_{\mathrm{o}}}$, and $\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{comp}})=\tfrac{(K-1)^{2}}{n_{\mathrm{c}}}\,\hat{q}(1-\hat{q}),\ \hat{q}=\tfrac{S_{\mathrm{c}}}{n_{\mathrm{c}}}.$ Then
@@
$\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{IVW}})=\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{ord}})\,\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{comp}})/(\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{ord}})+\widehat{\mathrm{Var}}(\widehat{A}_{\mathrm{comp}}))$. Substituting $q=\tfrac{A+K-2}{K-1}$ into Eq. (6) gives
@@
If $n_{\mathrm{c}}=0$, Eq. (9) reduces to $\widehat{A}_{\mathrm{ord}}$. If $n_{\mathrm{o}}=0$, it reduces to $\widehat{A}_{\mathrm{comp}}$. For $K=4$, one always has $\hat{q}\in[\tfrac{2}{3},1]$ when $A\in[0,1]$, which bounds the complementary-component variance. Moreover, a one-step Newton update around the truth shows that the IVW estimator with optimal fixed weights coincides with the ML estimator up to the first order [^2] A detailed proof is provided in Appendix D., and both are statistically efficient.
@@ 2.3 Theoretical analysis
We analyze the deviation of the complementary-label estimator $\widehat{A}_{\mathrm{comp}}$ under the uniform wrong-index assumption Eq. (1). Although $\widehat{A}_{\mathrm{comp}}$ is unbiased, its variance can be non-negligible at small sample sizes. We present two finite-sample deviation bounds: a distribution-free Hoeffding bound (Boucheron et al., 2003) and a data-dependent Bernstein bound (Maurer and Pontil, 2009) leveraging empirical variance. We then extend these to a fixed-weight mixture estimator.
2.3.1 Bounds for $\widehat{A}_{\mathrm{comp}}$
We derive a unified bound for $\widehat{A}_{\mathrm{comp}}$ by taking the minimum of the Hoeffding and empirical Bernstein inequalities. This formulation is both variance-free (via Hoeffding) and variance-adaptive (via Bernstein).The proof is shown in Appendix E.
@@ 2.3.1 Bounds for $\widehat{A}_{\mathrm{comp}}$
2.3.2 Bounds for $\hat{A}_{\mathrm{mix}}$
@@
We now state a Bernstein-type PAC bound for the mixed estimator that complements the union-bound inequalities established in Proposition 3; its proof is deferred to Appendix F.
@@
This bound tightens Hoeffding’s inequality by incorporating both the variance term $v$ and the (range-driven) linear term $c$. It is valid for any fixed choice of the weight $w\in[0,1]$, for instance if $w$ is set as a constant or estimated on an independent pilot set. If instead $w$ is estimated from the same evaluation data (e.g. the plug-in $\widehat{w}_{\mathrm{IVW}}$), the fixed-weight assumption is violated; in that case one must either revert to the union-bound bounds (Proposition 3), or apply standard remedies such as sample splitting (Chernozhukov et al., 2018; Wager and Athey, 2018)or a finite grid with a union bound (Boucheron et al., 2013; Shalev-Shwartz and Ben-David, 2014). For practical plug-in forms and further discussion, see Appendix G.